What minimum path length is needed to introduce a phase shift of 185° in light of wavelength 516 nm? (Give your answer in nm.)

What phase shift will that path difference you calculated in part (a) introduce in light of wavelength 750 nm? (Give you answer in degree.)
Answer #1

Solution)

a)

We know,

Phi = (2π/λ) * d * sinθ

Phi=phase difference

λ = wavelength

d * sinθ =path difference

or d * sinθ = (λ/2π) * φ

Given,

λ = 516 nm = 516 * 10^-9 m and phi =185o = 3.229 radian

Hence, d * sinθ = (516 * 10^-9/2π) * 3.229 =2.65* 10^-7 m

=========

b)

Wavelength, λ = 750 nm

φ = (2π/λ) * d * sinθ

Hence, φ = (2π/750 * 10^-9) * 2.65 *10^-7 = 2.22radian=127.19 degrees

=========